# Stochastic thermodynamic bounds on logical circuit operation

### v2

@@ III NOT Gate
$\chi_{\mathrm{p}}^{2}=\frac{\langle\tau_{\mathrm{p}}\rangle^{2}}{\langle\delta\tau_{\mathrm{p}}^{2}\rangle}\leq\frac{\langle Q\rangle}{2qV_{\mathrm{T}}}$ (2)
where brackets indicate an trajectory ensemble average, $\tau_{\mathrm{p}}$ is the first passage time to an output voltage passing the accuracy threshold $\alpha$, and $\delta x=(x-\langle x\rangle)$, and $Q$ is the the heat dissipated over a trajectory. Explicitly, $\tau_{\mathrm{p}}$ is calculated from an initial state corresponding to the steady-state of the inverter with the inverse input voltage, and the final absorbing boundary condition corresponds to a specific cumulative electronic current into the capacitor gate [50]. For a specific trajectory, the heat is given as
$Q=\ln\frac{P_{0}(x_{0})}{P_{f}(x_{f})}+\sum_{k=1}^{N}\ln\frac{W_{x_{k}x_{k-1}}}{W_{x_{k-1}x_{k}}}$ (3)
where the sum is over steps in the trajectory, and $x_{k}$ the state of the system at step $k$ [51]. The first term accounts for the specific starting and ending states of the trajectory, and the second term is equivalent to the general heat expression for an evolving distribution given above.
The thermodynamic uncertainty relation is a general result from stochastic thermodynamics, valid for any Markovian jump process. It states that the certainty in the first passage time, $\chi_{\mathrm{p}}$, is bounded from above by the heat dissipated over a trajectory, $Q$ [49]. The bound relates the minimum thermodynamic cost for a given desired certainty in the first passage time. While the bound was originally formulated for a system in steady state, here we apply it to a non-steady state system by including the additional first term associated with the change in system entropy in the trajectory heat expression above. On average, this system entropy term takes the form of a Kullback-Leibler divergence, which is by definition non-negative
$\langle\Delta s\rangle_{0}=\sum_{x_{0}}P_{0}(x_{0})\ln\frac{P_{0}(x_{0})}{P_{f}(x_{f})}=D(P_{0}\|P_{f})\geq 0$ (4)
and leads to a looser bound. Recent literature has suggested that tighter bounds can be obtained for transient, finite-time processes, but we do not apply them here [52, 53].
@@ IV Memory Device
Figure 2: Characterization of the memory device’s behavior and adherence to thermodynamic speed limits. (a) Probability of $V_{\mathrm{out}}$ as a function of $V_{\mathrm{d}}$, with crosses locating points of maximum probability at each $V_{\mathrm{d}}$ illustrating the onset of bistability. Example trajectories are shown for before and after the onset of bistability. (b) Mean time for a memory error as a function of the inverter relaxation time, controlled by increasing $V_{\mathrm{d}}$, as indicated on the top x-axis label, with the red dashed line indicating $\langle\tau_{\mathrm{err}}\rangle=\langle\tau_{\mathrm{p}}\rangle$. The shaded region is forbidden by the dissipation-time uncertainty relation.
@@
$\langle\tau_{\mathrm{err}}\rangle\geq\left(\beta\langle\dot{Q}\rangle\right)^{-1}$ (5)
where $\langle\dot{Q}\rangle$ is the average rate of heat dissipation in the steady-state. For this system, we work with an observation time over which the error rate is small, such that it is exponentially unlikely to see a back transition as required by the dissipation-time uncertainty relation. The region restricted by this lower bound is shown as the shaded area in Fig. 2b, indicating that the inequality is satisfied, but far from saturated. This result suggests that the memory stability time is not limited by heat dissipation.
Indeed this can be quantified using a nonequilibrium version of transition state theory [55, 56, 26]. Transition state theory bounds the rate of a transition between two metastable states using the stationary distribution $P_{\mathrm{ss}}(V_{\mathrm{out}})$ and an uncorrelated estimate of the time to cross a dividing surface between the two states. Equivalent to Kramers’ theory, the rate is estimated by the probability of a rare fluctuation, in this case a fluctuation of the output voltage of one of the NOT gates of the memory storage device being equal to $V_{\mathrm{d}}/2$. Taking the dividing surface to be $V_{\mathrm{out}}=V_{\mathrm{d}}/2$, and the typical time to relax from the top of the barrier as $\langle\tau_{\mathrm{p}}\rangle$, a nonequilibrium transition state theory estimate for $\langle\tau_{\mathrm{err}}\rangle$ is
@@ V Clock
Similarly, the dissipation-time TUR expressed in Eq. 5 can be adapted to yield a corresponding lower bound by replacing $\tau_{\mathrm{err}}$ for the memory device with $\tau_{\mathrm{c}}$ for the clock, which can be considered as a series of bit flips.
@@
$\chi_{\mathrm{c}}^{2}$ $\geq\frac{\langle\tau_{\mathrm{c}}\rangle}{\beta\langle\dot{Q}\rangle\langle\delta\tau_{\mathrm{c}}^{2}\rangle}$

### v3

@@ III NOT Gate
$\chi_{\mathrm{p}}^{2}=\frac{\langle\tau_{\mathrm{p}}\rangle^{2}}{\langle\delta\tau_{\mathrm{p}}^{2}\rangle}\leq\frac{\langle Q\rangle}{2qV_{\mathrm{T}}}+1$ (2)
where brackets indicate an trajectory ensemble average, $\tau_{\mathrm{p}}$ is the first passage time to an output voltage passing the accuracy threshold $\alpha$, and $\delta x=(x-\langle x\rangle)$, and $Q$ is the the heat dissipated over a trajectory. Explicitly, $\tau_{\mathrm{p}}$ is calculated from an initial state $X=0$ (or $X=1$), and the final absorbing boundary condition corresponds to a specific cumulative electronic current into (or out of) the capacitor gate [50]. For a specific trajectory, the heat is given as
$Q=\sum_{k=1}^{N}\ln\frac{W_{x_{k}x_{k-1}}}{W_{x_{k-1}x_{k}}}$ (3)
where the sum is over steps in the trajectory, and $x_{k}$ the state of the system at step $k$ [51].
The thermodynamic uncertainty relation is a general result from stochastic thermodynamics, valid for any Markovian jump process. It states that the certainty in the first passage time, $\chi_{\mathrm{p}}$, is bounded from above by the heat dissipated over a trajectory, $Q$ [52]. The bound relates the minimum thermodynamic cost for a given desired certainty in the first passage time. While the bound was originally formulated for a system in steady state, here we apply a finite-time version [49].
@@ IV Memory Device
Figure 2: Characterization of the memory device’s behavior and adherence to thermodynamic speed limits. (a) Probability of $V_{\mathrm{out}}$ as a function of $V_{\mathrm{d}}$, with crosses locating points of maximum probability at each $V_{\mathrm{d}}$ illustrating the onset of bistability. Example trajectories are shown for before and after the onset of bistability. (b) Mean time for a memory error as a function of the inverter relaxation time, controlled by increasing $V_{\mathrm{d}}$, as indicated on the top x-axis label.
@@
The efficiency of this memory device can be quantified using a nonequilibrium version of transition state theory [54, 55, 26]. Transition state theory bounds the rate of a transition between two metastable states using the stationary distribution $P_{\mathrm{ss}}(V_{\mathrm{out}})$ and an uncorrelated estimate of the time to cross a dividing surface between the two states. Equivalent to Kramers’ theory, the rate is estimated by the probability of a rare fluctuation, in this case a fluctuation of the output voltage of one of the NOT gates of the memory storage device being equal to $V_{\mathrm{d}}/2$. Taking the dividing surface to be $V_{\mathrm{out}}=V_{\mathrm{d}}/2$, and the typical time to relax from the top of the barrier as $\langle\tau_{\mathrm{p}}\rangle$, a nonequilibrium transition state theory estimate for $\langle\tau_{\mathrm{err}}\rangle$ is
@@ V Clock
Similarly, the dissipation-time uncertainty relation [24] relates the rate of heat dissipation with the mean time to complete a process. It can be applied for the clock cycle time to yield a lower bound,
@@
$\chi_{\mathrm{c}}^{2}$ $\geq\frac{\langle\tau_{\mathrm{c}}\rangle}{\beta\langle\dot{Q}\rangle\langle\delta\tau_{\mathrm{c}}^{2}\rangle}\,,$